How many kilograms of zinc that is 33% pure should be mixed with 12 kilograms of alloy that is 45% pure to make an alloy that is 39% pure zinc?
Answer by Paul(988)

0.33x+0.45(12)=0.39(x+12)
0.33x+5.4=0.39x+4.68
-0.06x=-0.72
x=12

Hence, about 12 solution must be added.
Paul.